The body of mass 5kg is suspended by an inextensible string from a nail 0 and is pulled by a horizontal force F until the angle of inclination of the string to the vertical is 60 degrees. Calculate the value of F and the tension in the string( g as 10m/s
1
Expert's answer
2020-10-12T07:50:04-0400



"m\overrightarrow{g}+\overrightarrow{F}+\overrightarrow{T}=\overrightarrow{0}"

"\begin{cases}
 -F+Tsin\alpha=0 \\
 -mg+Tcos\alpha=0
\end{cases}"

"\begin{cases}
 T=\frac{mg}{cos\alpha} \\
 F=Tsin\alpha
\end{cases}"

"\begin{cases}
 T=\frac{mg}{cos\alpha} \\
 F=mgtan\alpha
\end{cases}"

"T=\frac{5 \cdot 10}{0.5}=100" (N)

"F=5\cdot10\cdot1.73=86.5" (N).

Answer: 86.5 N, 100 N.